# A Little Depth Goes a Long Way: The Expressive Power of Log-Depth Transformers

### v2

@@ Appendix D Graph Connectivity Proof
We will prove this for directed graphs, as an undirected edge between two vertices can be equivalently represented as two directed edges between those vertices. Let $G$ be a directed graph over $n$ vertices. Let $A\in\{0,1\}^{n\times n}$ be $G$’s adjacency matrix: for $i,j\in\{1,\ldots,n\}$, $A_{i,j}$ is $1$ if $G$ has an edge from $i$ to $j$, and 00 otherwise.
@@
We first argue that $B_{\lceil\log n\rceil}(i,j)=1$ if and only if $G$ has a path from $i$ to $j$. Clearly, there is such a path if and only if there is a “simple path” of length at most $n$ from $i$ to $j$. To this end, we argue by induction over $\ell$ that $B_{\ell}(i,j)=1$ if an only if $G$ has a path of length at most $2^{\ell}$ from $i$ to $j$. For the base case of $\ell=0$, by construction, $B_{0}(i,j)=1$ if and only if either $i=j$ (which we treat as a path of length 00) or $A_{i,j}=1$ (i.e., there is a direct edge from $i$ to $j$). Thus, $B_{\ell}(i,j)=1$ if and only if there is a path of length at most $12^{0}=$ 1 from $i$ to $j$. Now suppose the claim holds for $B_{\ell}(i,j)$. By construction, $C_{\ell+1}(i,k,j)=1$ if and only if $B_{\ell}(i,k)=B_{\ell}(k,j)=1$, which by induction means there are paths of length at most $2^{\ell}$ from $i$ to $k$ and from $k$ to $j$, which in turn implies that there is a path of length at most $2\cdot 2^{\ell}=2^{\ell+1}$ from $i$ to $j$ (through $k$). Furthermore, note conversely that if there is a path of length at most $2^{\ell+1}$ from $i$ to $j$, then there must exist a “mid-point” $k$ in this path such that there are paths of length at most $2^{\ell}$ from $i$ to $k$ and from $k$ to $j$, i.e., $C_{\ell+1}(i,k,j)=1$ for some $k$. This is precisely what the definition of $B_{\ell+1}(i,j)$ captures: it is $1$ if and only if there exists a $k$ such that $C_{\ell+1}(i,k,j)=1$, which, as argued above, holds if and only if there is a path of length at most $2^{\ell+1}$ from $i$ to $j$. This completes the inductive step.
@@
The transformer starts off by using layer 1 to store $1/N,n,n^{2},s,$ and $t$ in the residual stream at every position, as follows. The layer uses one head with uniform attention and with value $1$ only at the first token (recall that the position embedding is assumed to separate 1 from other positions). This head computes $1/N$ and the layer adds $\psi(1/N)$ to the residual stream. Note that the input tokens in the first set of $n^{2}$ positions, namely 00 and $1$, are distinct from tokens in the rest of the input. The layer, at every position, uses a second head with uniform attention, and with value $1$ at tokens in $\{0,1\}$ and value 00 at all other tokens. This head computes $n^{2}/N$. The layer now adds $\psi(n^{2}/N,1/N)$, where $\psi(a,b)$ is defined as the (unnormalized) vector $\langle a,b,-a,-b\rangle$. When these coordinates are later read from the residual stream via masked pre-norm, they will get normalized and one would obtain $\phi(n^{2}/N,1/N)=\phi(n^{2})$. Thus, future layers will have access to $\phi(n^{2})$ through the residual stream. The layer similarly uses three additional heads to compute $n^{3}/N$, $s/N$, and $t/N$. From the latter two values, it computes $\psi(s/N,1/N)$ and $\psi(t/N,1/N)$ and adds them to the residual stream; as discussed above, these can be read in future layers as $\phi(s/N,1/N)=\phi(s)$ and $\phi(t/N,1/N)=\phi(t)$. Finally, the layer computes $\psi(n^{3}/N,n^{2}/N)$ and adds it to the residual stream. Again, this will be available to future layers as $\phi(n^{3}/N,n^{2}/N)=\phi(n)$.
@@
The transformer uses the next 15 layers to compute and store in the residual stream the semantic “coordinates” of each of the first $n^{2}+n^{3}$ token position as follows. For each of the first $n^{2}$ positions $p=in+j$ with $1\leq p\leq n^{2}$, it uses Lemma 1 (7 layers) with $a_{i}$ set to $p$ and $m$ set $n$ in order to add $\phi(i)$ and $\phi(j)$ to the residual stream at position $p$. In parallel, for each of the next $n^{3}$ positions $p=n^{2}+(in^{2}+kn+j)$ with $n^{2}+1\leq p\leq n^{2}+n^{3}$, it uses Lemma 1 with $a_{i}$ set to $p$ and $m$ set $n$ in order to add $\phi((i+1)n+k)$ and $\phi(j)$ to the residual stream. It then uses the lemma again (7 more layers), this time with $a_{i}$ set to $(i+1)n+k$ and $m$ again set to $n$, to add $\phi(i+1)$ and $\phi(k)$ to the residual stream. Lastly, it uses Lemma 7 applied to $\phi(i+1)$ to add $\phi(i)$ to the residual stream.
@@
Layer 17 of the transformer computes the predicate $B_{0}(i,j)$ at the first $n^{2}$ token positions as follows. At position $p=in+j$, it uses Lemma 8 to compute $\mathbb{I}(\phi(A(i,j)=\phi(1))$ and $\mathbb{I}(\phi(i)=\phi(j))$; note that $\phi(A(i,j))$, $\phi(i)$, and $\phi(j)$ are available in the residual stream at position $p$. It then uses a feedforward layer to output $1$ if both of these are $1$, and output 00 otherwise. This is precisely the intended value of $B_{0}(i,j)$. The sublayer then adds $B_{0}(i,j)$ to the residual stream. The layer also adds to the residual stream the value $1$, which will be used to initialize the boolean that controls layer alternation in the repeated layers as discussed next.
@@
The next set of layers alternates between computing the $C_{\ell}$ and the $B_{\ell}$ predicates for $\ell\in\{1,\ldots,\lceil\log n\rceil\}$. To implement this, each position $i$ at layer updates in the residual stream the value of a single boolean $r$ computed as follows. $r$ is initially set to $1$ at layer 8. Each repeating layer retrieves $r$ from the residual stream and adds $1-r$ to the same coordinate in the residual stream. The net effect is that the value of $r$ alternates between $1$ and 00 at the repeating layers. The transformer uses this to alternate between the computation of the $C_{\ell}$ and the $B_{\ell}$ predicates.
@@
For $\ell\in\{1,\ldots,\lceil\log n\rceil\}$, layer $(2\ell-1)+8$ of the transformer computes the predicate $C_{\ell}(i,k,j)$ at the set of $n^{3}$ (padding) positions $p=n^{2}+in^{2}+kn+j$, as follows. It uses two heads, one with query $\langle\phi(i),\phi(k)\rangle$ and the other with query $\langle\phi(k),\phi(j)\rangle$. The keys in the first $n^{2}$ positions $q=i^{\prime}n+j^{\prime}$ are set to $\langle\phi(i^{\prime}),\phi(j^{\prime})\rangle$, and the values are set to $B_{\ell-1}(i^{\prime},j^{\prime})$. The two heads thus attend solely to positions with coordinates $(i,k)$ and $(k,j)$, respectively, and retrieve boolean values $B_{\ell-1}(i,k)$ and $B_{\ell-1}(k,j)$, respectively, stored there in the previous layer. The layer then uses Lemma 8 to compute $\mathbb{I}(B_{\ell-1}(i,k)=1)$ and $\mathbb{I}(B_{\ell-1}(k,j)=1)$, and uses a feedforward layer to output $1$ if both of these checks pass, and output 00 otherwise. This is precisely the intended value of $C_{\ell}(i,k,j)$. If $\ell>1$, the layer replaces the value $C_{\ell-1}(i,k,j)$ stored previously in the residual stream with the new boolean value $C_{\ell}(i,k,j)$ by adding $C_{\ell}(i,k,j)-C_{\ell-1}(i,k,j)$ to the same coordinates of the residual stream. If $\ell=1$, it simply adds $C_{\ell}(i,k,j)$ to the residual stream.
@@
For $\ell\in\{1,\ldots,\lceil\log n\rceil\}$, layer $82\ell+$ 8 computes the predicate $B_{\ell}(i,j)$ at the first $n^{2}$ positions $p=in+j$, as follows. It uses a head with query $\langle\phi(i),\phi(j)\rangle$. The keys in the second set of $n^{3}$ positions $q=n^{2}+i^{\prime}n^{2}+k^{\prime}n+j^{\prime}$ are set to $\langle\phi(i^{\prime}),\phi(j^{\prime})\rangle$ (recall that $\phi(i^{\prime})$ and $\phi(j^{\prime})$ are available in the residual stream at $q$) and the corresponding values are set to the boolean $C_{\ell}(i^{\prime},k^{\prime},j^{\prime})$, stored previously in the residual stream. The head thus attends uniformly to the $n$ padding positions that have coordinates $(i,k^{\prime},j)$ for various choices of $k^{\prime}$. It computes the average of their values, which equals $h=\frac{1}{n}\sum_{k^{\prime}=1}^{n}C_{\ell}(i,k^{\prime},j)$ as well as $1/(2n)$ using an additional head. We observe that $h\geq 1/n$ if there exists a $k^{\prime}$ such that $C_{\ell}(i,k^{\prime},j)=1$, and $h=0$ otherwise. These conditions correspond precisely to $B_{\ell}(i,j)$ being $1$ and 00, respectively. We compute $h-1/(2n)$ and store it in the residual stream. Similar to the proof of Lemma 8, the feedforward layer reads $\sigma=\mathrm{sgn}(h-1/(2n))$, computes $z=(1+\mathrm{ReLU}(\sigma))/2$, and writes $z$ to the residual stream. The value $z$ is precisely the desired $B_{\ell}(i,j)$ as $\sigma$ is $1$ when $h\geq 1/n$ and 00 when $h=0$. As in Lemma 8, the intermediate value $h-1/(2n)$ written to the residual stream can be recomputed and reset in the next layer. As before, the transformer replaces the value $B_{\ell-1}(i,j)$ stored previously in the residual stream with the newly computed value $B_{\ell}(i,j)$ by adding $\psi(B_{\ell}(i,j)-B_{\ell-1}(i,j))$ to the stream at the same coordinates.
@@
Finally, in layer $182\lceil\log n\rceil+$ 18, the final token uses a head that attends with query $\langle\phi(s),\phi(t)\rangle$ corresponding to the source and target nodes $s$ and $t$ mentioned in the input; recall that $\phi(s)$ and $\phi(t)$ are available in the residual stream. The keys in the first $n^{2}$ positions $p=in+j$ are, as before, set to $\langle\phi(i),\phi(j)\rangle$, and the values are set to $B_{\lceil\log n\rceil}(i,j)$ retrieved from the residual stream. The head thus attends solely to the position with coordinates $(s,t)$, and retrieves and outputs the value $B_{\lceil\log n\rceil}(s,t)$. This value, as argued earlier, is $1$ if and only if $G$ has a path from $s$ to $t$. ∎
@@ Appendix E Proofs for Width Scaling and Chain of Thought Claims
By assumption, we can construct an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit family in which the transformer weights for sequence length $n$ are hardcoded as constants. Next, we can apply standard arguments [Merrill et al., 2022, Merrill and Sabharwal, 2023a, b] to show that the self-attention and feedforward sublayers can both be simulated by constant-depth threshold circuits, and the size remains polynomial (though a larger polynomial). Thus, any function computable by a constant-depth, polynomial-width transformer is in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. ∎
@@
The high-level idea is that a polynomial-size circuit can enumerate all possible $O(\log n)$-length chains of thought. Then, in parallel for each chain of thought, we construct a threshold circuit that simulates a transformer [Merrill and Sabharwal, 2023a] on the input concatenated with the chain of thought, outputting the transformer’s next token. We then select the chain of thought in which all simulated outputs match the correct next token and output its final answer. The overall circuit has constant depth, polynomial size, and can be shown to be $\mathsf{L}$-uniform. Thus, any function computable by a transformer with $O(\log n)$ chain of thought is in $\mathsf{TC}^{0}$. ∎
@@ Appendix F Experimental Details
In early experiments, we found that learning from long $A_{5}$ sequences directly was infeasible for our transformer models. We hypothesize this was because, unless earlier tokens are predicted correctly, later tokens contribute significant noise to the gradient. In order to make the learning problem feasible, we follow a curriculum training process, first training on $A_{5}$ sequences of length 2, then length 4, and continuing up to some fixed maximum power $2^{i}$. We can then measure the maximum $n^{*}\leq 2^{i}$ such that the model achieves strong validation accuracy, as mentioned in Section 7.
@@
All depth experiments used a fixed width of 512. For historical reasons, we have slightly different numbers of runs for different experimental conditions, and some of the runs use different batch sizes (64 and 128). We originally ran a single sweep of depths and widths with 5 runs for depths 6, 12, 18, and 24, each using a batch size of 64 and maximum depth of $1282^{i}=$ 128. Seeking to clarify the trend between these original data points, we launched 3 additional runs at depths 9, 15, 18, and 21 using a batch size of 128, which anecdotally sped up training without harming final performance. We also observed that the original depth 24 runs were at the ceiling $n^{*}=128$, so we launched 3 additional depth-24 runs with a batch size of 128 and $5122^{i}=$ 512 (we also used this larger sequence length for all other runs in the second set). In total, this made:
@@
All width experiments used a fixed depth of 6. We launched 5 runs at widths 128, 258, 512, 1024 with the same hyperparameters, each using a batch size of 64 and $1282^{i}=$ 128.

### v3

@@ Appendix D Graph Connectivity Proof
We will prove this for directed graphs, as an undirected edge between two vertices can be equivalently represented as two directed edges between those vertices. Let $G$ be a directed graph over $n$ vertices. Let $A\in\{0,1\}^{n\times n}$ be $G$’s adjacency matrix: for $i,j\in\{1,\ldots,n\}$, $A_{i,j}$ is $1$ if $G$ has an edge from $i$ to $j$, and 0 otherwise.
@@
We first argue that $B_{\lceil\log n\rceil}(i,j)=1$ if and only if $G$ has a path from $i$ to $j$. Clearly, there is such a path if and only if there is a “simple path” of length at most $n$ from $i$ to $j$. To this end, we argue by induction over $\ell$ that $B_{\ell}(i,j)=1$ if an only if $G$ has a path of length at most $2^{\ell}$ from $i$ to $j$. For the base case of $\ell=0$, by construction, $B_{0}(i,j)=1$ if and only if either $i=j$ (which we treat as a path of length 0) or $A_{i,j}=1$ (i.e., there is a direct edge from $i$ to $j$). Thus, $B_{\ell}(i,j)=1$ if and only if there is a path of length at most $2^{0}=1$ from $i$ to $j$. Now suppose the claim holds for $B_{\ell}(i,j)$. By construction, $C_{\ell+1}(i,k,j)=1$ if and only if $B_{\ell}(i,k)=B_{\ell}(k,j)=1$, which by induction means there are paths of length at most $2^{\ell}$ from $i$ to $k$ and from $k$ to $j$, which in turn implies that there is a path of length at most $2\cdot 2^{\ell}=2^{\ell+1}$ from $i$ to $j$ (through $k$). Furthermore, note conversely that if there is a path of length at most $2^{\ell+1}$ from $i$ to $j$, then there must exist a “mid-point” $k$ in this path such that there are paths of length at most $2^{\ell}$ from $i$ to $k$ and from $k$ to $j$, i.e., $C_{\ell+1}(i,k,j)=1$ for some $k$. This is precisely what the definition of $B_{\ell+1}(i,j)$ captures: it is $1$ if and only if there exists a $k$ such that $C_{\ell+1}(i,k,j)=1$, which, as argued above, holds if and only if there is a path of length at most $2^{\ell+1}$ from $i$ to $j$. This completes the inductive step.
@@
The transformer starts off by using layer 1 to store $1/N,n,n^{2},s,$ and $t$ in the residual stream at every position, as follows. The layer uses one head with uniform attention and with value $1$ only at the first token (recall that the position embedding is assumed to separate 1 from other positions). This head computes $1/N$ and the layer adds $\psi(1/N)$ to the residual stream. Note that the input tokens in the first set of $n^{2}$ positions, namely 0 and $1$, are distinct from tokens in the rest of the input. The layer, at every position, uses a second head with uniform attention, and with value $1$ at tokens in $\{0,1\}$ and value 0 at all other tokens. This head computes $n^{2}/N$. The layer now adds $\psi(n^{2}/N,1/N)$, where $\psi(a,b)$ is defined as the (unnormalized) vector $\langle a,b,-a,-b\rangle$. When these coordinates are later read from the residual stream via masked pre-norm, they will get normalized and one would obtain $\phi(n^{2}/N,1/N)=\phi(n^{2})$. Thus, future layers will have access to $\phi(n^{2})$ through the residual stream. The layer similarly uses three additional heads to compute $n^{3}/N$, $s/N$, and $t/N$. From the latter two values, it computes $\psi(s/N,1/N)$ and $\psi(t/N,1/N)$ and adds them to the residual stream; as discussed above, these can be read in future layers as $\phi(s/N,1/N)=\phi(s)$ and $\phi(t/N,1/N)=\phi(t)$. Finally, the layer computes $\psi(n^{3}/N,n^{2}/N)$ and adds it to the residual stream. Again, this will be available to future layers as $\phi(n^{3}/N,n^{2}/N)=\phi(n)$.
@@
The transformer uses the next 15 layers to compute and store in the residual stream the semantic “coordinates” of each of the first $n^{2}+n^{3}$ token position as follows. For each of the first $n^{2}$ positions $p=in+j$ with $1\leq p\leq n^{2}$, it uses Lemma˜1 (7 layers) with $a_{i}$ set to $p$ and $m$ set $n$ in order to add $\phi(i)$ and $\phi(j)$ to the residual stream at position $p$. In parallel, for each of the next $n^{3}$ positions $p=n^{2}+(in^{2}+kn+j)$ with $n^{2}+1\leq p\leq n^{2}+n^{3}$, it uses Lemma˜1 with $a_{i}$ set to $p$ and $m$ set $n$ in order to add $\phi((i+1)n+k)$ and $\phi(j)$ to the residual stream. It then uses the lemma again (7 more layers), this time with $a_{i}$ set to $(i+1)n+k$ and $m$ again set to $n$, to add $\phi(i+1)$ and $\phi(k)$ to the residual stream. Lastly, it uses Lemma˜7 applied to $\phi(i+1)$ to add $\phi(i)$ to the residual stream.
@@
Layer 17 of the transformer computes the predicate $B_{0}(i,j)$ at the first $n^{2}$ token positions as follows. At position $p=in+j$, it uses Lemma˜8 to compute $\mathbb{I}(\phi(A(i,j)=\phi(1))$ and $\mathbb{I}(\phi(i)=\phi(j))$; note that $\phi(A(i,j))$, $\phi(i)$, and $\phi(j)$ are available in the residual stream at position $p$. It then uses a feedforward layer to output $1$ if both of these are $1$, and output 0 otherwise. This is precisely the intended value of $B_{0}(i,j)$. The sublayer then adds $B_{0}(i,j)$ to the residual stream. The layer also adds to the residual stream the value $1$, which will be used to initialize the boolean that controls layer alternation in the repeated layers as discussed next.
@@
The next set of layers alternates between computing the $C_{\ell}$ and the $B_{\ell}$ predicates for $\ell\in\{1,\ldots,\lceil\log n\rceil\}$. To implement this, each position $i$ at layer updates in the residual stream the value of a single boolean $r$ computed as follows. $r$ is initially set to $1$ at layer 8. Each repeating layer retrieves $r$ from the residual stream and adds 1r1-r to the same coordinate in the residual stream. The net effect is that the value of $r$ alternates between $1$ and 0 at the repeating layers. The transformer uses this to alternate between the computation of the $C_{\ell}$ and the $B_{\ell}$ predicates.
@@
For $\ell\in\{1,\ldots,\lceil\log n\rceil\}$, layer $(2\ell-1)+8$ of the transformer computes the predicate $C_{\ell}(i,k,j)$ at the set of $n^{3}$ (padding) positions $p=n^{2}+in^{2}+kn+j$, as follows. It uses two heads, one with query $\langle\phi(i),\phi(k)\rangle$ and the other with query $\langle\phi(k),\phi(j)\rangle$. The keys in the first $n^{2}$ positions $q=i^{\prime}n+j^{\prime}$ are set to $\langle\phi(i^{\prime}),\phi(j^{\prime})\rangle$, and the values are set to $B_{\ell-1}(i^{\prime},j^{\prime})$. The two heads thus attend solely to positions with coordinates $(i,k)$ and $(k,j)$, respectively, and retrieve boolean values $B_{\ell-1}(i,k)$ and $B_{\ell-1}(k,j)$, respectively, stored there in the previous layer. The layer then uses Lemma˜8 to compute $\mathbb{I}(B_{\ell-1}(i,k)=1)$ and $\mathbb{I}(B_{\ell-1}(k,j)=1)$, and uses a feedforward layer to output $1$ if both of these checks pass, and output 0 otherwise. This is precisely the intended value of $C_{\ell}(i,k,j)$. If $\ell>1$, the layer replaces the value $C_{\ell-1}(i,k,j)$ stored previously in the residual stream with the new boolean value $C_{\ell}(i,k,j)$ by adding $C_{\ell}(i,k,j)-C_{\ell-1}(i,k,j)$ to the same coordinates of the residual stream. If $\ell=1$, it simply adds $C_{\ell}(i,k,j)$ to the residual stream.
@@
For $\ell\in\{1,\ldots,\lceil\log n\rceil\}$, layer $2\ell+8$ computes the predicate $B_{\ell}(i,j)$ at the first $n^{2}$ positions $p=in+j$, as follows. It uses a head with query $\langle\phi(i),\phi(j)\rangle$. The keys in the second set of $n^{3}$ positions $q=n^{2}+i^{\prime}n^{2}+k^{\prime}n+j^{\prime}$ are set to $\langle\phi(i^{\prime}),\phi(j^{\prime})\rangle$ (recall that $\phi(i^{\prime})$ and $\phi(j^{\prime})$ are available in the residual stream at $q$) and the corresponding values are set to the boolean $C_{\ell}(i^{\prime},k^{\prime},j^{\prime})$, stored previously in the residual stream. The head thus attends uniformly to the $n$ padding positions that have coordinates $(i,k^{\prime},j)$ for various choices of $k^{\prime}$. It computes the average of their values, which equals $h=\frac{1}{n}\sum_{k^{\prime}=1}^{n}C_{\ell}(i,k^{\prime},j)$ as well as $1/(2n)$ using an additional head. We observe that $h\geq 1/n$ if there exists a $k^{\prime}$ such that $C_{\ell}(i,k^{\prime},j)=1$, and $h=0$ otherwise. These conditions correspond precisely to $B_{\ell}(i,j)$ being $1$ and 0, respectively. We compute h1/(2n)h-1/(2n) and store it in the residual stream. Similar to the proof of Lemma˜8, the feedforward layer reads $\sigma=\mathrm{sgn}(h-1/(2n))$, computes $z=(1+\mathrm{ReLU}(\sigma))/2$, and writes $z$ to the residual stream. The value $z$ is precisely the desired $B_{\ell}(i,j)$ as $\sigma$ is $1$ when $h\geq 1/n$ and 0 when $h=0$. As in Lemma˜8, the intermediate value h1/(2n)h-1/(2n) written to the residual stream can be recomputed and reset in the next layer. As before, the transformer replaces the value $B_{\ell-1}(i,j)$ stored previously in the residual stream with the newly computed value $B_{\ell}(i,j)$ by adding $\psi(B_{\ell}(i,j)-B_{\ell-1}(i,j))$ to the stream at the same coordinates.
@@
Finally, in layer $2\lceil\log n\rceil+18$, the final token uses a head that attends with query $\langle\phi(s),\phi(t)\rangle$ corresponding to the source and target nodes $s$ and $t$ mentioned in the input; recall that $\phi(s)$ and $\phi(t)$ are available in the residual stream. The keys in the first $n^{2}$ positions $p=in+j$ are, as before, set to $\langle\phi(i),\phi(j)\rangle$, and the values are set to $B_{\lceil\log n\rceil}(i,j)$ retrieved from the residual stream. The head thus attends solely to the position with coordinates $(s,t)$, and retrieves and outputs the value $B_{\lceil\log n\rceil}(s,t)$. This value, as argued earlier, is $1$ if and only if $G$ has a path from $s$ to $t$. ∎
@@ Appendix E Proofs for Width Scaling and Chain of Thought Claims
By assumption, we can construct an $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit family in which the transformer weights for sequence length $n$ are hardcoded as constants. Next, we can apply standard arguments [merrill2022SatAttnTC0, merrill-sabharwal-2023-parallelism, merrill2023logic] to show that the self-attention and feedforward sublayers can both be simulated by constant-depth threshold circuits, and the size remains polynomial (though a larger polynomial). Thus, any function computable by a constant-depth, polynomial-width transformer is in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. ∎
@@
The high-level idea is that a polynomial-size circuit can enumerate all possible $O(\log n)$-length chains of thought. Then, in parallel for each chain of thought, we construct a threshold circuit that simulates a transformer [merrill-sabharwal-2023-parallelism] on the input concatenated with the chain of thought, outputting the transformer’s next token. We then select the chain of thought in which all simulated outputs match the correct next token and output its final answer. The overall circuit has constant depth, polynomial size, and can be shown to be $\mathsf{L}$-uniform. Thus, any function computable by a transformer with $O(\log n)$ chain of thought is in $\mathsf{TC}^{0}$. ∎
@@ Appendix F Experimental Details
In early experiments, we found that learning from long $A_{5}$ sequences directly was infeasible for our transformer models. We hypothesize this was because, unless earlier tokens are predicted correctly, later tokens contribute significant noise to the gradient. In order to make the learning problem feasible, we follow a curriculum training process, first training on $A_{5}$ sequences of length 2, then length 4, and continuing up to some fixed maximum power $2^{i}$. We can then measure the maximum $n^{*}\leq 2^{i}$ such that the model achieves strong validation accuracy, as mentioned in Section˜7.
@@
All depth experiments used a fixed width of 512. For historical reasons, we have slightly different numbers of runs for different experimental conditions, and some of the runs use different batch sizes (64 and 128). We originally ran a single sweep of depths and widths with 5 runs for depths 6, 12, 18, and 24, each using a batch size of 64 and maximum depth of $2^{i}=128$. Seeking to clarify the trend between these original data points, we launched 3 additional runs at depths 9, 15, 18, and 21 using a batch size of 128, which anecdotally sped up training without harming final performance. We also observed that the original depth 24 runs were at the ceiling $n^{*}=128$, so we launched 3 additional depth-24 runs with a batch size of 128 and $2^{i}=512$ (we also used this larger sequence length for all other runs in the second set). In total, this made:
@@
All width experiments used a fixed depth of 6. We launched 5 runs at widths 128, 258, 512, 1024 with the same hyperparameters, each using a batch size of 64 and $2^{i}=128$.